OpenStudy (cookiegirl):

Please help i will give you a medal! :D Which expressions are equivalent to the expression 20 ÷ (4 – 2)? Choose exactly two answers that are correct. A. 8 + 12 ÷ 2 B. 14 – 8 ÷ 2 C. 2 • (7 – 4) D. 16 – 14 ÷ 2 + 1

OpenStudy (welshfella):

what does 20 / (4 - 2) equal? work out what is in the parentheses first

OpenStudy (cookiegirl):

10?

OpenStudy (welshfella):

yes now work out the 4 options and see which of these also equal 10 remember the order of operations.

OpenStudy (cookiegirl):

umm ok thanks....

OpenStudy (welshfella):

for example A:- do the division first

OpenStudy (welshfella):

8 + 12 / 2 = 8 + 6

OpenStudy (cookiegirl):

its 14

OpenStudy (welshfella):

- so A is not = to 10.

OpenStudy (welshfella):

yes so you can cross off A

OpenStudy (cookiegirl):

and B is 3 so its not B either

OpenStudy (cookiegirl):

so its C and D?

OpenStudy (welshfella):

No B is not 3 DO the DIVISION first.

OpenStudy (cookiegirl):

I did....

OpenStudy (welshfella):

i mad a mistake like that earlier on!

OpenStudy (welshfella):

14 - 8 / 2 = 14 - 4

OpenStudy (cookiegirl):

so its B and D?....

OpenStudy (welshfella):

yes

OpenStudy (cookiegirl):

Thanks for your help...

OpenStudy (welshfella):

yw